OpenStudy (hulagirlmb):

My teacher is out of the office today and I was wondering about a test question and why I got it wrong. Can somebody explain it to me? I'll put it in the replies.

OpenStudy (hulagirlmb):

\[2\sqrt{5}+3\sqrt{5}-\sqrt{5}\]

OpenStudy (hulagirlmb):

I put in my answer as \[5\sqrt{5}\] and the computer said it was wrong and the answer was \[4\sqrt{5}\]

OpenStudy (hulagirlmb):

@iGreen

OpenStudy (anonymous):

The answer is indeed 4√5, and i'll explain why.

OpenStudy (hulagirlmb):

ok..

OpenStudy (anonymous):

\[2√5+3√5−√5~\rightarrow ~2√5+3√5~\rightarrow~ 1√5−√5~\rightarrow~4√5\]

OpenStudy (hulagirlmb):

it's just a bunch of question marks

OpenStudy (anonymous):

Refresh your page.